YOU ARE SENT TO THE STORE TO BUY TOMATOES AND BROCCOLI. THREE POUNDS OF TOMATOES AND FOUR POUNDS OF BROCCOLI WILL COST $12. TWO POUNDS OF TOMATOES AND TWO POUNDS OF BROCCOLI WILL COST $8. WHAT IS THE INDIVIDUAL COST OF THE TOMATOES AND BROCCOLI?
Answer by stanbon(48568)

THREE POUNDS OF TOMATOES AND FOUR POUNDS OF BROCCOLI WILL COST $12. TWO POUNDS OF TOMATOES AND TWO POUNDS OF BROCCOLI WILL COST $8. WHAT IS THE INDIVIDUAL COST OF THE TOMATOES AND BROCCOLI?
------------
3T + 4B = 12
2T + 2B = 8
-------------
Multiply 2nd by 2 to get:
4T + 4B = 16
-----------------
Subtract that from 1st to get:
-T = -4
T = 4 dollars (price of a pound of tomatoes)
---------------
Substitute to solve for B:
3*4 + 4B = 12
4B = 0
B = 0 (price of a pound of broccoli)
============
Cheers,
Stan H.